# Searching for GEMS: The Occurrence of Giant Planets orbiting M-dwarfs within 100 pc

### v1

@@ 6.3 Future Work
This work to search a 100 pc M-dwarf sample is a pilot study to prepare for the analysis of a 200 pc M-dwarf sample, which represents a further eight-fold increase in sample size. Work beyond the 200 pc sample will aim to produce occurrence rates for giant exoplanets orbiting all M-dwarfs observed by TESS. The larger sample will also aid in adopting more sophisticated occurrence rate estimation techniques such as an HBM approach with astrophysical stellar dependent priors or the ABC method.
Future work will include the additional observation and characterization of the new GEMS candidate identified in this work, TIC 146712644. This search also yielded 267 non-noise false positives at the manual vetting stage, including 142 circularized EBs, 19 eccentric EBs, 69 contact binaries, and two other transiting systems with periods beyond the scope of this study. Further work on these FPs will utilize spectroscopic observations, which coupled with TESS photometry, and ground-based photometric follow-up will allow us to confirm the false positive nature of these objects. These observations of double-lined EBs, which make up the majority of the FP sample, provide some of the best and most direct measurements of masses, radii, and limb darkening coefficients, which are especially useful in a regime where asteroseismology is not effective (Henry & Jao, 2024). These model-independent measurements of the M-dwarfs can then be used to compare and constrain theoretical models of low mass stars near the hydrogen-burning limit, a regime where models still show significant discrepancies from observations (Baroch et al., 2018), and will help to enable a more accurate characterization of the planets orbiting them (Benedict et al., 2016). The sample of FPs will also allow us to help form statistics of close-in binaries around M-dwarf primaries, helping to inform the statistics of close companions to M-dwarf primaries, which will help reconcile the differences seen between planet occurrence rates derived from radial velocity surveys with transit surveys (Moe & Kratter, 2021).
In addition to direct outgrowths of this study, this work emphasizes the need for future work across several different areas. 14(a) demonstrates the need for more precise constraints on GEMS with orbital periods longer than 10 days. Such constraints might be enabled by the upcoming PLAnetary Transits and Oscillations of stars mission (PLATO; Rauer et al. (2025)), which includes M-dwarfs in its target list ($M_{*}<0.5~\mathrm{M_{\odot}}$) and a nominal mission life of >4>4 years. These estimates would help to determine whether an ice line occurrence rate turnover similar to that observed by Fernandes et al. (2019) in FGK-star giant planets exists for GEMS as well. Furthermore, 14(b) notes the lack of occurrence rate data for super-Neptunes orbiting M-dwarfs.
@@ 7 Summary
As part of the Searching for GEMS survey, we performed a pilot study of a distance-limited (<< 100 pc) sample of M-dwarfs observed by TESS using our new TESS-miner package. Each of our candidates was subjected to further exhaustive vetting, pixel-checks and transit fitting, following which we were left with 23 candidates. Of these, 7 were existing planets, 3 had previously been identified as astrophysical FPs by past research, and 12 were identified as astrophysical false positives based on spectroscopic observations, leaving 1 new candidate requiring later spectroscopic vetting. Based on this, we calculated occurrence rates for GEMS around early-type ($0.067\%\pm 0.047\%$), mid-type ($0.139\%\pm 0.069\%$ or a maximum of $0.139^{+0.072}_{-0.062}\%$ including the new planet candidates), and late-type M-dwarfs ($0.032\%\pm 0.032\%$). We also calculated an overall GEMS occurrence rate across the complete M-dwarf sample to be $0.065^{+0.025}_{-0.027}\%$, or a maximum of $0.065^{+0.026}_{0.026}\%$ including new planet candidates. Our estimates on occurrence of giant planets across the M-dwarf spectral type, when combined with literature estimates for FGK stars, enabled us to fit an exponential relation between giant planet occurrence rates and stellar mass, which confirms the rarity of GEMS due to the lower average mass protoplanetary disks on average around M-dwarfs.
@@
Resources supporting this work were provided by the (i) NASA High-End Computing Program through the NASA Center for Climate Simulation (NCCS) at Goddard Space Flight Center (ii) Pennsylvania State University’s Institute for Computational and Data Sciences’ (ICDS) Roar supercomputer, and (iii) the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology. This content is solely the responsibility of the authors and does not necessarily represent the views of the NCCS, ICDS, or Caltech. These results are based on observations obtained with the Habitable-zone Planet Finder Spectrograph on the HET. We acknowledge support from NSF grants AST-1006676, AST-1126413, AST-1310885, AST-1310875, ATI 2009889, ATI-2009982, AST-2108512, AST-2108801 and the NASA Astrobiology Institute (NNA09DA76A) in the pursuit of precision radial velocities in the NIR. The HPF team also acknowledges support from the Heising-Simons Foundation via grant 2017-0494. The Hobby-Eberly Telescope is a joint project of the University of Texas at Austin, the Pennsylvania State University, Ludwig-Maximilians-Universität München, and Georg-August Universität Gottingen. The HET is named in honor of its principal benefactors, William P. Hobby and Robert E. Eberly. The HET collaboration acknowledges the support and resources from the Texas Advanced Computing Center. We thank the Resident astronomers and Telescope Operators at the HET for the skillful execution of our observations with HPF. We would like to acknowledge that the HET is built on Indigenous land. Moreover, we would like to acknowledge and pay our respects to the Carrizo & Comecrudo, Coahuiltecan, Caddo, Tonkawa, Comanche, Lipan Apache, Alabama-Coushatta, Kickapoo, Tigua Pueblo, and all the American Indian and Indigenous Peoples and communities who have been or have become a part of these lands and territories in Texas, here on Turtle Island. WIYN is a joint facility of the University of Wisconsin-Madison, Indiana University, NSF’s NOIRLab, the Pennsylvania State University, Purdue University, University of California-Irvine, and the University of Missouri. This research was carried out, in part, at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors are honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham. Data presented herein were obtained at the WIYN Observatory from telescope time allocated to NN-EXPLORE through the scientific partnership of NASA, the NSF, and NOIRLab (program IDs 2023B-438370 and 2024A-103024, both with PI Kanodia). This paper uses WINERED data gathered with the 6.5 meter Magellan Telescope located at Las Campanas Observatory, Chile. We are grateful to Shogo Otsubo, Yuki Saragaku, and Tomomi Takeuchi (Kyoto-Sangyo University), and to Noriyuki Matsunaga (University of Tokyo) of the WINERED team, as well as the staff of Las Campanas Observatory, for their support during the WINERED observations. WINERED was developed by the University of Tokyo and the Laboratory of Infrared High-resolution Spectroscopy, Kyoto Sangyo University, under the financial support of KAKENHI (Nos. 16684001, 20340042, and 21840052) and the MEXT Supported Program for the Strategic Research Foundation at Private Universities (Nos. S0801061 and S1411028). The observing run in 2023 June was partly supported by KAKENHI (grant No. 18H01248) and JSPS Bilateral Program Number JPJSBP120239909. Andrew McWilliam thanks and acknowledges receipt of a Carnegie Venture Grant, kindly provided by the Carnegie Institution for Science, in order to purchase equipment required to adapt, install and support WINERED on the Magellan Clay telescope. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Some of the data presented in this paper were obtained from MAST at STScI. Support for MAST for non-HST data is provided by the NASA Office of Space Science via grant NNX09AF08G and by other grants and contracts. This work includes data collected by the TESS mission, which are publicly available from MAST. Funding for the TESS mission is provided by the NASA Science Mission directorate. This research made use of the (i) NASA Exoplanet Archive, which is operated by Caltech, under contract with NASA under the Exoplanet Exploration Program, (ii) SIMBAD database, operated at CDS, Strasbourg, France, (iii) NASA’s Astrophysics Data System Bibliographic Services, and (iv) data from 2MASS, a joint project of the University of Massachusetts and IPAC at Caltech, funded by NASA and the NSF. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France, and NASA’s Astrophysics Data System Bibliographic Services. This research has made use of the Exoplanet Follow-up Observation Program website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. CIC acknowledges support by an appointment to the NASA Postdoctoral Program at the Goddard Space Flight Center, administered by ORAU through a contract with NASA. We thank Thomas Barclay for sharing the latest Roman GBTDS pointings. We thank Corey Beard for assistance in developing and testing sub-modules related to differences between odd and even transits. Finally, we thank the anonymous referee for their careful review and thoughtful suggestions, which have significantly improved the clarity and quality of this work.

### v3

@@ 6.3 Future Work
This search within a 100 pc M-dwarf sample is a pilot study to prepare for the analysis of a 200 pc M-dwarf sample, which represents a further eight-fold increase in sample size. Work beyond the 200 pc sample will aim to produce occurrence rates for giant exoplanets orbiting all M-dwarfs observed by TESS. The larger sample will also aid in adopting more sophisticated occurrence rate estimation techniques such as an HBM approach with astrophysical stellar dependent priors or the ABC method.
This search also yielded 267 non-noise false positives at the manual vetting stage, including 142 circularized EBs, 19 eccentric EBs, 69 contact binaries, and two other transiting systems with periods beyond the scope of this study. Further work on these FPs will utilize spectroscopic observations, which coupled with TESS photometry, and ground-based photometric follow-up will allow us to confirm the false positive nature of these objects. These observations of double-lined EBs, which make up the majority of the FP sample, provide some of the best and most direct measurements of masses, radii, and limb darkening coefficients, which are especially useful in a regime where asteroseismology is not effective (henry_character_2024). These model-independent measurements of the M-dwarfs can then be used to compare and constrain theoretical models of low mass stars near the hydrogen-burning limit, a regime where models still show significant discrepancies from observations (Baroch2018), and will help to enable a more accurate characterization of the planets orbiting them (benedict_solar_2016). The sample of FPs will also allow us to help form statistics of close-in binaries around M-dwarf primaries, helping to inform the statistics of close companions to M-dwarf primaries, which will help reconcile the differences seen between planet occurrence rates derived from radial velocity surveys with transit surveys (Moe2021).
In addition to direct outgrowths of this study, this work emphasizes the need for future work across several different areas. 14(a) demonstrates the need for more precise constraints on GEMS with orbital periods longer than 10 days. Such constraints might be enabled by the upcoming PLAnetary Transits and Oscillations of stars mission (PLATO; Rauer2025), which includes M-dwarfs in its target list ($M_{*}<0.5~\mathrm{M_{\odot}}$) and a nominal mission life of >4>4 years. These estimates would help to determine whether an ice line occurrence rate turnover similar to that observed by fernandes_hints_2019 in FGK-star giant planets exists for GEMS as well. Furthermore, 14(b) notes the lack of occurrence rate data for super-Neptunes orbiting M-dwarfs.
@@ 7 Summary
As part of the Searching for GEMS survey, we performed a pilot study of a distance-limited (<< 100 pc) sample of M-dwarfs observed by TESS using our new TESS-miner package. Each of our candidates was subjected to further exhaustive vetting, pixel-checks and transit fitting, following which we were left with 23 candidates. Of these, 7 were existing planets, 3 had previously been identified as astrophysical FPs by past research, and 13 were identified as astrophysical false positives based on spectroscopic observations. Based on this, we calculated occurrence rates for GEMS around early-type ($0.118\%\pm 0.068\%$), mid-type ($0.153\%\pm 0.069\%$), and late-type M-dwarfs ($0.036\%\pm 0.024\%$). We also calculated an overall GEMS occurrence rate across the complete M-dwarf sample to be $0.068\%\pm 0.024\%$. Our estimates on occurrence of giant planets across the M-dwarf spectral type, when combined with literature estimates for FGK stars, enabled us to fit an exponential relation between giant planet occurrence rates and stellar mass, which confirms the rarity of GEMS due to the lower average mass protoplanetary disks on average around M-dwarfs.
@@
Resources supporting this work were provided by the (i) NASA High-End Computing Program through the NASA Center for Climate Simulation (NCCS) at Goddard Space Flight Center (ii) Pennsylvania State University’s Institute for Computational and Data Sciences’ (ICDS) Roar supercomputer, and (iii) the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology. This content is solely the responsibility of the authors and does not necessarily represent the views of the NCCS, ICDS, or Caltech. These results are based on observations obtained with the Habitable-zone Planet Finder Spectrograph on the HET. We acknowledge support from NSF grants AST-1006676, AST-1126413, AST-1310885, AST-1310875, ATI 2009889, ATI-2009982, AST-2108512, AST-2108801 and the NASA Astrobiology Institute (NNA09DA76A) in the pursuit of precision radial velocities in the NIR. The HPF team also acknowledges support from the Heising-Simons Foundation via grant 2017-0494. The Hobby-Eberly Telescope is a joint project of the University of Texas at Austin, the Pennsylvania State University, Ludwig-Maximilians-Universität München, and Georg-August Universität Gottingen. The HET is named in honor of its principal benefactors, William P. Hobby and Robert E. Eberly. The HET collaboration acknowledges the support and resources from the Texas Advanced Computing Center. We thank the Resident astronomers and Telescope Operators at the HET for the skillful execution of our observations with HPF. We would like to acknowledge that the HET is built on Indigenous land. Moreover, we would like to acknowledge and pay our respects to the Carrizo & Comecrudo, Coahuiltecan, Caddo, Tonkawa, Comanche, Lipan Apache, Alabama-Coushatta, Kickapoo, Tigua Pueblo, and all the American Indian and Indigenous Peoples and communities who have been or have become a part of these lands and territories in Texas, here on Turtle Island. WIYN is a joint facility of the University of Wisconsin-Madison, Indiana University, NSF’s NOIRLab, the Pennsylvania State University, Purdue University, University of California-Irvine, and the University of Missouri. This research was carried out, in part, at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors are honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham. Data presented herein were obtained at the WIYN Observatory from telescope time allocated to NN-EXPLORE through the scientific partnership of NASA, the NSF, and NOIRLab (program IDs 2023B-438370 and 2024A-103024, both with PI Kanodia). This paper uses WINERED data gathered with the 6.5 meter Magellan Telescope located at Las Campanas Observatory, Chile. We are grateful to Shogo Otsubo, Yuki Saragaku, and Tomomi Takeuchi (Kyoto-Sangyo University), and to Noriyuki Matsunaga (University of Tokyo) of the WINERED team, as well as the staff of Las Campanas Observatory, for their support during the WINERED observations. WINERED was developed by the University of Tokyo and the Laboratory of Infrared High-resolution Spectroscopy, Kyoto Sangyo University, under the financial support of KAKENHI (Nos. 16684001, 20340042, and 21840052) and the MEXT Supported Program for the Strategic Research Foundation at Private Universities (Nos. S0801061 and S1411028). The observing run in 2023 June was partly supported by KAKENHI (grant No. 18H01248) and JSPS Bilateral Program Number JPJSBP120239909. Andrew McWilliam thanks and acknowledges receipt of a Carnegie Venture Grant, kindly provided by the Carnegie Institution for Science, in order to purchase equipment required to adapt, install and support WINERED on the Magellan Clay telescope. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Some of the data presented in this paper were obtained from MAST at STScI. Support for MAST for non-HST data is provided by the NASA Office of Space Science via grant NNX09AF08G and by other grants and contracts. This work includes data collected by the TESS mission, which are publicly available from MAST. Funding for the TESS mission is provided by the NASA Science Mission directorate. This research made use of the (i) NASA Exoplanet Archive, which is operated by Caltech, under contract with NASA under the Exoplanet Exploration Program, (ii) SIMBAD database, operated at CDS, Strasbourg, France, (iii) NASA’s Astrophysics Data System Bibliographic Services, and (iv) data from 2MASS, a joint project of the University of Massachusetts and IPAC at Caltech, funded by NASA and the NSF. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France, and NASA’s Astrophysics Data System Bibliographic Services. This research has made use of the Exoplanet Follow-up Observation Program website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. CIC acknowledges support by an appointment to the NASA Postdoctoral Program at the Goddard Space Flight Center, administered by ORAU through a contract with NASA. We thank Thomas Barclay for sharing the latest Roman GBTDS pointings. We thank Corey Beard for assistance in developing and testing sub-modules related to differences between odd and even transits. Finally, we thank the anonymous referee for their careful review and thoughtful suggestions, which have significantly improved the clarity and quality of this work. Figure 15 shows an example of each of the 7 types of FPs identified by eye in Stage 2 of our analysis, along with an example of a light curve which survived Stage 2.
Figure 15: Example detrended and phase-folded light curves vetted manually in Section 4. All figures are direct outputs of the TESS-miner package. a. shows an example of instrumental noise being falsely caught by the BLS search as a planet candidate. b. shows an example of the stellar rotation signal doing the same. c. shows an example of an EB FP, where the BLS search retrieved twice the period, therefore invalidating the OddEvenFlag. d. shows an example of an EB FP, where the BLS search retrieved half the period. e. shows an example of an eccentric EB FP, where the phase-folded light curve reveals an occultation offset from the expectation for a circular orbit. f. shows an example of a contact binary EB FP. g. shows an example of an otherwise acceptable planet candidate where the true period is outside out scope of interest (either $P<1$ day or $P>10$ days). Finally, h. is an example of an acceptable candidate.
Figure 16 shows an example of a TGLC pixel grid where light from the target star is uncontaminated by other sources (a) and an example of a pixel grid where this is not the case (b). In the latter case, a candidate is removed from consideration. This pixel-pixel check was performed during Stage 2 of analysis.
(a) The pixel-pixel normalised light curves for TIC 139811797, Sector 12, one of our new planet candidates. The transit signal is visible in the pixels centred on the position of the target and do not appear to originate from any other star. For this reason, this object passed our pixel-pixel check.
(b) The pixel-pixel normalised light curves for TIC 376843143, Sector 25. The transit signal is visible in the pixels centred on TIC 376843132 and not the target star. For this reason, this objected did not pass our pixel-pixel check.
Figure 16: Top: an example of a pixel-pixel check pass, and Bottom: an example of a pixel-pixel check failure. These figures are generated by the tglc package.
Table 7 compiles each TOI within our input sample not explicitly noted in the main text of this work where $1.0~\mathrm{days}<P_{\text{TOI}}<10.0$ days and $0.7<R_{\text{TOI}}/R_{J}<1.5$. We explain briefly in which stage each object was eliminated and why.
Table 7: All TOI within our input sample and search criteria with which we have have not otherwise explicitly dealt in this work.
TIC TOI Details
146846569 734.01 This object survived until Stage 2a. We removed it as a likely EB due to its V-shaped transit and evidence of a secondary eclipse, particularly in Sector 36.
67646988 1779.01 This object also survived until the manual vetting stage, when we removed it due to its fitted $R_{p}/R_{*}=0.35$ and V-shaped transit.
151728428 5850.01 This object was classified as a Poor Fit during Stage 0. Several factors contributed to this classification. The object was only observed in one sector within the TESS PM and EM1 (Sector 54). Within that sector, an excess of noise led to a significant fraction of the light curve being cut from the analysis. Only one transit occurred within the remaining data, and consequently the AntiTransitBLSFlag1 and the RadiusFlag were both triggered due to poor fitting. In addition, this object was retired as an SB1 by the TOI working group in May, 2024.